OpenStudy (anonymous):

One leg of a 45°-45°-90° triangle is 3√2 . Find the length of the hypotenuse. Question 13 options: 4 5 6 8

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):



OpenStudy (dtan5457):

if \[x=3\sqrt{2}\]

OpenStudy (dtan5457):

then you need to do \[3\sqrt{2}\times \sqrt{2}\]

OpenStudy (dtan5457):

which is 3\[\sqrt{4}\]

OpenStudy (dtan5457):

radical 4=2, 3 times 2=6<<ur answer

OpenStudy (anonymous):

wow u deserve a hug

OpenStudy (dtan5457):

lol. those questions before i'm sorry they are wrong for some reason. i seriously think they are right though, something is wrong with the answer key or something.

OpenStudy (anonymous):

its okay . i emailed my teacher . i think you right also

OpenStudy (anonymous):

but can you look at the other question i tag you in or can i put all the questions on the same chat or you want me to do it seperately ?